OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6)

OpenStudy (hoblos):

parallel lines have same slopes so the slope is -3 y-6 = -3(x+1) y-6 = -3x -3 y = -3x +3